MsBrains (ms-brains):

Solve, using the substitution method. j + k = 3 j – k = 7 A.The solution is (8, 1) B.The solution is (5, –2) C.There is no solution. D.There are an infinite number of solutions.

OpenStudy (caozeyuan):

if j-k=7, what is j?

MsBrains (ms-brains):

If it was substituted it would be A.8 - 1 = 7 And the other one on top would be 5 - 2 = 3

OpenStudy (caozeyuan):

but x and y are same throut

OpenStudy (caozeyuan):

so if j-k=7 j=7+k, and j+k=3 so 7+2k=3, now what is k

OpenStudy (serenity74):

I messed up j + k = 3 j - k = 7 cancel out the ks you end up with 2j = 10 j = 5 then plug back the 5 into one of the equations and solve for k

MsBrains (ms-brains):

Wait, Okay. I don't understand.

OpenStudy (serenity74):

first equation (5) + k = 3 k = -5 +3 = -2 or second equation (5) - k = 7 -k = 7-5 = 2 BUT k is negative and it can't be negative so you bring the negative sign onto the other side and you end up with k = -2

MsBrains (ms-brains):

The answer has to be C

OpenStudy (caozeyuan):

no it is b

OpenStudy (serenity74):

No its B according to my calculations. Then what's the answer and please show me the work.

MsBrains (ms-brains):

How can the solution be (5, –2) ?

OpenStudy (serenity74):

I just showed you all the math up there.....

OpenStudy (serenity74):

@DanJS what do you think?

OpenStudy (danjs):

been remembering that the the product of the x coefficeients + product of the y coefficients, will be zero, perpendicular lines but that is not what you do here..

OpenStudy (anonymous):

if you need any help learning the topic more use the link here https://www.khanacademy.org/math/cc-eighth-grade-math/cc-8th-systems-topic/cc-8th-systems-with-substitution/v/the-substitution-method it will help you learn more on it

OpenStudy (danjs):

j + k = 3 j - k = 7 there will be some solution where they intersect at point (j,k) from looking at it, ( k + 7) + k = 3 2k = -4 k = -2 and then j will be 5

OpenStudy (danjs):

that is the unique solution, the two lines cross at the point (j,k) = (5, -2)

OpenStudy (danjs):

try it in the problem 5 - 2 = 3 and 5+2 = 7

OpenStudy (serenity74):

See I was right! thank you @DanJS